# Cosmic Cartography II: completing galaxy catalogs for gravitational-wave cosmology

### v1

@@ Abstract
The dark siren method exploits the complementarity between gravitational-wave binary coalescence signals and galaxy catalogs originating from the same regions of space. However, all galaxy catalogs are incomplete, i.e. they only include a subset of all galaxies, typically being biased towards the bright end of the luminosity distribution. This sub-selection systematically affects the dark siren inference of the Hubble constant $H_{0}$, so a completeness relation has to be introduced that accounts for the missing objects. In the literature it is standard to assume that the missing galaxies are uniformly distributed across the sky and that the galaxy magnitude distribution is known. In this work we develop a novel method which improves upon these assumptions and reconstructs the underlying true galaxy field, respecting the spatial correlation of galaxies on large scales. In our method the true magnitude distribution of galaxies is inferred alongside the spatial galaxy distribution. Our method results in an improved three-dimensional prior in redshift and sky position for the host galaxy of a GW event, which is expected to make the resulting $H_{0}$ posterior more robust. Building on our previous work, we make a number of improvements, and validate our method on simulated data based on the Millennium simulation.
@@ 1.1 Current state of the field
Gravitational waves (GWs) provide an independent means of probing the expansion history of the Universe. This is particularly interesting given the latest discrepancies in the cosmological parameters, both in the current expansion rate of the Universe, $H_{0}$ [1, 2], and the evolving equation of state of dark energy [3, 4]. The extent to which GWs can help resolve these growing tensions is now a central question in the field [5, 6, 7, 8, 9, 10, 11]. LIGO-Virgo-KAGRA (LVK) GW data [12, 13, 14, 15, 16] have been used to measure the Hubble constant through the multi-messenger event GW170817 [17, 18], and also through the analysis of the GW transient catalog GWTC-3 [19], resulting in $H_{0}=68^{+12}_{-8}\,\,{\rm km\,s^{-1}\,Mpc^{-1}}$ [20]. For the Hubble constant measurement it is a large advantage to have precise redshift information, that can either originate from a direct measurement of the host galaxy redshift [17, 18, 21, 22, 23, 24, 25, 26] (i.e. the bright siren method), from a statistical measurement using a galaxy catalog [27, 28, 5, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42] (i.e. the dark siren method), or the redshifted source-frame mass distribution [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58] (i.e. the spectral siren method). [^1] Additionally, there has been a vivid recent interest in the development of GW-galaxy cross-correlation techniques that forecast a $H_{0}$ measurement with percent-level accuracy [59, 60] for future, simulated data.
@@
Dark siren codes for $H_{0}$ inference [35, 38, 7] now combine and leverage both galaxy and source-frame mass distributions into cosmological measurements, although the galaxy catalog information currently provides little constraining power [61, 20, 35, 38]. This is usually attributed to (i) the incompleteness of the available galaxy catalogs (i.e. faint galaxies are only seen in the near-by Universe) as well as (ii) most GW signals are not well-localized in the sky or luminosity distance. In case of (i) little redshift information for the GW is provided, and hence, the constraint on $H_{0}$ is weak. Even if the catalog is complete, a large GW sky position (point (ii)) leads to averages of the galaxy redshift prior over large volumes. Since the Universe is homogeneous on large scales, this case is asymptotically equivalent (in the limit of no sky position information) to using solely the marginalized redshift distribution of galaxies. Thus, poorly localized GW events yield weaker constraints on cosmological parameters when compared to well-localized events.
@@
The latest LVK cosmological analyses [20] leverage GW data in combination with the set of galaxy catalogs called glade+ [62, 63]. This set of catalogs is essentially empty in the K-band at $z\sim 0.3$. Since many GW signals occur at higher equivalent redshifts (equivalent assuming a reference cosmology to convert the GW luminosity distance to a redshift), the catalog incompleteness poses a bottleneck. To formulate a completeness correction (i.e. what fraction of galaxies is observed for a given redshift), a galaxy magnitude distribution is required, the parameters of which are fixed in all standard codes, although this distribution is not known perfectly. Additionally, the aforementioned codes [35, 38, 7] implement the completion correction assuming a uniform-in-comoving-distance distribution for the GW host. While it is acknowledged that the missing galaxies are not uniformly scattered, little progress has been made to complete galaxy catalogs disposing of this unphysical assumption, for exceptions see [33] and [64, 65, 66]).
@@
Our works solves the above problems. We address the incompleteness, by reconstructing the full galaxy field given a magnitude-limited catalog, under the hypothesis that the galaxy field respects a given two-point statistic (accounting for spatial correlation). We also marginalize over the galaxy magnitude distribution; it is jointly inferred along with the galaxy field. This leads to a more informative galaxy redshift-sky position prior for GW events that implies a more informative dark siren $H_{0}$ measurement, schematically illustrated in Sec. 1.2. While this work demonstrates the reconstruction on simulated data it prepares the method for its application to glade+.
@@ 1.2 Representative toy model
For simplicity, we linearize the relationship between redshift and luminosity distance, i.e. $d_{L}=c\,z/H_{0}$. All analyzed GW events are localized in the same sky pixel (i.e. element of the sky) that contains $\mathcal{O}(10^{3})$ galaxies (detected and non-detected). To illustrate the impact of the galaxy completeness we then vary the number of observed galaxies included in the catalog. Each galaxy has a redshift uncertainty of $\sigma_{z}=0.01$. From the observed galaxy catalog one can build the line-of-sight redshift prior, the sum of the individual galaxy redshift probability density distributions. [^2] As pointed out in [39] this treatment is approximate, since one should consider a hierarchical model for the galaxy distribution. Currently however, this treatment is standard in all inference codes. Note that because the catalog is incomplete, the line-of-sight redshift prior has two contributions: the in-catalog part (informed by the observed galaxies) and out-of-catalog part (with the standard assumption that galaxies are distributed uniformly in comoving volume). We construct the in-catalog part as a sum of Gaussian distributions with their respective redshift error that modulate the volumetric $z^{2}$ prior. The out-of-catalog part is simply modeled as a $z^{2}$ distribution. If we define the sigmoid
@@
Figure 1: Toy model: schematic impact of the galaxy catalog completeness. Left panel: The redshift prior for one fiducial sky pixel, where different colors indicate different completeness. The catalogs transition to the homogeneous distribution at different redshifts: while the catalog of 10 galaxies (orange) reverts to the homogeneous distribution already at $z\sim 0.2$, the catalog of 800 galaxies (blue) only transitions at $z\sim 1$. This has important consequences on the $H_{0}$ posterior as the right panel illustrates: while the most incomplete catalog provides an almost uninformative $H_{0}$ posterior, the more complete catalogs with 200 and 800 galaxies, respectively, lead to a more informative $H_{0}$ measurement.
@@ 1.3 Novelty
Let us here clearly draw up the differences between our previous method of [65] and the present work. In our previous reconstruction we have estimated the galaxy number count per three-dimensional voxel. We had assumed a simplified setup, namely a cartesian box, a parametrized form for the galaxy magnitude distribution, and a local bias prescription linking dark matter (DM) density contrast and mean galaxy rate. For the DM model itself, we assumed a log-normal field with a two-point statistic that was estimated but of a parametrized form. Here, we go significantly further and
@@ 1.4 Paper structure
Figure 3: Schematic overview of the hierarchical inference problem. Starting from the true catalog, one identifies possible galaxies (step Object detection). We indicate the apparent magnitude (cf. Eq. (2.8)) of each galaxy with its color, darker (blue) signifying an apparent brighter galaxy. In a second step (step Inference galaxy properties) the parameters of each galaxy are inferred, independently for each object. Finally, the inference of the individual galaxy properties is improved by considering a joint prior on the overall distribution of galaxies (step Solve inverse problem). In this step, we start from the magnitude-limited catalog with the noisy galaxy properties and build possible configurations of the true catalog that are compatible with the observations.
@@ 2 Forward-modeling of galaxy catalog data
${n_{{\rm c},\hat{I}}}$ Observed number count in bin $\hat{I}$ (measured quantities)
@@
Schematically, the likelihood can be written as ($\mathbb{P}(\{\mathcal{D}\}|\Lambda)$, with $\{\mathcal{D}\}\coloneqq\{\mathcal{D}_{c}\}_{c\in C}$ the observed catalog where $c$ is a label running over all galaxies in the observed catalog $C$. The raw catalog data comes in the form of fluxes in different bands, or spectra from which the galaxy properties (redshift, composition, sky position, star formation rates, etc.) are deduced. We will refer to these individual properties as $\theta$. As such, $\theta$ is a $\mathcal{O}(3-20)$-dimensional vector. One possible way to characterize each galaxy is through posterior samples that approximate $\theta\sim\mathbb{P}(\theta|\mathcal{D})\,,$ the individual posterior of the galaxy properties. The likelihood of observing the catalog data $\{\mathcal{D}\}$ given the hyperparameters $\Lambda$ can be computed through [^3] In this expression we omit terms relating to the selection effect and that depend on the overall rate. This equation is for illustrating purposes only: this approach would be computationally prohibitive and will not be used in the later analysis.
@@
marginalizing the joint prior of the true catalog, ($\mathbb{P}(\{\theta_{c}\}_{c\in C}|\Lambda)$, over the individual galaxy properties. While future computational acceleration might allow the evaluation of the above integral, it is not the aim of this work to marginalize over the properties of $\mathcal{O}(10^{6-8})$ galaxies.
Therefore, we choose to represent the catalog data as the galaxies histogrammed over the sky, the redshift and the apparent magnitude. Since the redshift is uncertain, we bin each galaxy by its maximum-likelihood value in redshift. That is, our observable is a number count in each three-dimensional bin, which we denote as ${n_{{\rm c},\hat{I}}}$, where $\hat{I}$ is the index in redshift, sky position and magnitude, and the hat indicating that the galaxies are binned via their noisy maximum-likelihood properties. In general, the maximum-likelihood values do not correspond to their true values, an effect accounted for in our modeling, as outlined below.
@@
($\mathbb{P}(\{\mathcal{D}\}|\Lambda)\approx\mathbb{P}(\{{n_{{\rm c},\hat{I}}}\}|\Lambda)\,.$ (2.2)
Two assumptions underlie the above approximation, namely (i) the probability of $\mathbb{P}(\mathcal{D}|\Lambda)$ is constant in each bin and, (ii) each galaxy data can be fully described by the bin in which it falls (here its maximum-likelihood value for its properties). While the first assumption can be satisfied by using a fine grid over the galaxy properties, the second assumption is stronger. If we assume the $\theta$ uncertainties to be known, this assumes the individual galaxy likelihood to be independent of other latent variables not included in the histogram, e.g. the time of observation, the type of galaxy, etc.
@@
One can consider the problem we attempt to solve as a hierarchical inference problem where the individually obtained galaxy properties (which were independently inferred from the other galaxy data) are improved through the formulation of a joint prior of galaxy properties of position and magnitude, accounting for the three-dimensional spatial correlation and magnitude distribution. This is illustrated by Fig. 3.
@@ 2.2 Dark matter model
In order to reconstruct the full galaxy field, we choose to model the mean galaxy number in each three-dimensional voxel as a local function of the DM density contrast. [^4] As discussed later in Sec. 2.3, this is an approximation, as the real galaxy formation process is a complex function of the halo size, galaxy history, environment etc. To produce the latter (spatially correlated) field in spherical coordinates, we now describe the log-normal field [67] that models the DM density contrast. Since we produce DM realizations on a cartesian grid, but need to evaluate its density on a spherical grid, we detail our discretization scheme and how we interpolate from cartesian to spherical coordinates in App. B.
Following the methodology of [65] based on [67], we generate a log-normal field with a given power spectrum on a cartesian grid. The power spectrum is computed using a phenomenological model that we introduce below in Eq. (2.5), the parameters of which we fit jointly with the other remaining hyperparameters such as the DM density contrast. The log-normal field is generated by applying a transformation to a Gaussian random field, given by [67, 68]
@@
where we denote the Gaussian random field as $F_{\rm G}$, and $\sigma_{\text{G}}^{2}$ is the variance of the Gaussian random field (over many field realizations). The Gaussian field is first constructed by sampling from a multivariate normal distribution with a covariance matrix derived from the desired power spectrum, exploiting an analytical connection between power spectrum of the log-normal field and its underlying Gaussian random field [67]. The log-normal transformation of Eq. (2.4) ensures that the resulting field is bounded by $-1$ from below, as required for the DM density contrast. Within this formalism, we do not take into account the redshift dependence of the power spectrum.
@@
$\mathcal{P}_{\text{DM}}(\vec{k})=P_{A,\text{DM}}\,\frac{||\vec{k}||^{P_{n,1}}}{P_{\xi}+\left(\frac{||\vec{k}||}{P_{k,\text{eq}}}\right)^{P_{n,2}}}\,,$ (2.5)
where $\vec{k}$ is the (three-dimensional) wave vector and ||.||||.||| | . | | is the Euclidean norm. We show two examples of the power spectrum and their corresponding log-normal fields in two dimensions in Fig. 4. Two reasons made us choose a phenomenological model for the DM power spectrum: (i) our previous work [65] found little impact of using a power spectrum that depends on cosmological parameters and (ii) it is faster to calculate.

### v3

@@ Abstract
The dark siren method exploits the complementarity between gravitational-wave binary coalescence signals and galaxy catalogs originating from the same regions of space. However, all galaxy catalogs are incomplete, i.e. they only include a subset of all galaxies, typically being biased towards the bright end of the luminosity distribution. This sub-selection systematically affects the dark siren inference of the Hubble constant $H_{0}$, so a completeness relation has to be introduced that accounts for the missing objects. In the literature it is standard to assume that the missing galaxies are uniformly distributed across the sky and that the galaxy magnitude distribution is known. In this work we develop a novel method which improves upon these assumptions and reconstructs the underlying true galaxy field, respecting the spatial correlation of galaxies on large scales. In our method the true magnitude distribution of galaxies is inferred alongside the spatial galaxy distribution. Our method results in an improved three-dimensional prior in redshift and sky position for the host galaxy of a GW event, which is expected to make the resulting $H_{0}$ posterior more robust. Building on our previous work, we make a number of improvements, and validate our method on simulated data based on the Millennium simulation. The inference results can be reproduced through our publicly available code base  light.
@@ 1.1 Current state of the field
Gravitational waves (GWs) provide an independent means of probing the expansion history of the Universe. This is particularly interesting given the latest discrepancies in the cosmological parameters, both in the current expansion rate of the Universe, $H_{0}$ [Planck:2018vyg, Riess:2021jrx], and the evolving equation of state of dark energy [DESI:2024mwx, DESI:2025zgx]. The extent to which GWs can help resolve these growing tensions is now a central question in the field [LIGOScientific:2018gmd, Borhanian:2020vyr, Borghi:2023opd, Hanselman:2024hqy, Wang:2024zfv, Chen:2024gdn, Tagliazucchi:2025ofb]. LIGO-Virgo-KAGRA (LVK) GW data [det1-aligo2015, det2-aLIGO:2020wna, det3-Tse:2019wcy, det4-VIRGO:2014yos, det5-Virgo:2019juy] have been used to measure the Hubble constant through the multi-messenger event GW170817 [LIGOScientific:2017vwq, LIGOScientific:2017adf], and also through the analysis of the GW transient catalog GWTC-3 [KAGRA:2021vkt], resulting in $H_{0}=68^{+12}_{-8}\,\,{\rm km\,s^{-1}\,Mpc^{-1}}$ [LIGOScientific:2021aug]. For the Hubble constant measurement it is a large advantage to have precise redshift information, that can either originate from a direct measurement of the host galaxy redshift [LIGOScientific:2017vwq, LIGOScientific:2017adf, Howlett:2019mdh, Chen:2020dyt, Chen:2023dgw, Palmese:2023beh, Mancarella:2024qle, Salvarese:2024jpq] (i.e. the bright siren method), from a statistical measurement using a galaxy catalog [Schutz:1986gp, DelPozzo:2011vcw, LIGOScientific:2018gmd, Gray:2019ksv, DES:2020nay, Gray:2021sew, Palmese:2021mjm, Finke:2021aom, Turski:2023lxq, Mastrogiovanni:2023emh, DESI:2023fij, Alfradique:2023giv, Gray:2023wgj, Gair:2022zsa, Perna:2024lod, Beirnaert:2025wcx, Naveed:2025kgk] (i.e. the dark siren method), or the redshifted source-frame mass distribution [Taylor:2011fs, Taylor:2012db, Farr:2019twy, Mastrogiovanni:2021wsd, Mancarella:2021ecn, Leyde:2022orh, Ezquiaga:2022zkx, Pierra:2023deu, Leyde:2023iof, Bom:2024afj, Farah:2024xub, MaganaHernandez:2024uty, Mali:2024wpq, Agarwal:2024hld, Li:2024rmi, Tong:2025xvd] (i.e. the spectral siren method).
@@
Dark siren codes for $H_{0}$ inference [Mastrogiovanni:2023emh, Gray:2023wgj, Borghi:2023opd] now combine and leverage both galaxy and source-frame mass distributions into cosmological measurements, although the galaxy catalog information currently provides little constraining power [LIGOScientific:2019zcs, LIGOScientific:2021aug, Mastrogiovanni:2023emh, Gray:2023wgj]. This is usually attributed to (i) the incompleteness of the available galaxy catalogs (i.e. faint galaxies are only seen in the near-by Universe) as well as (ii) most GW signals are not well-localized in the sky or luminosity distance. In case of (i) little redshift information for the GW is provided, and hence, the constraint on $H_{0}$ is weak. Even if the catalog is complete, a large GW sky position (point (ii)) leads to averages of the galaxy redshift prior over large volumes. Since the Universe is homogeneous on large scales, this case is asymptotically equivalent (in the limit of no sky position information) to using solely the marginalized redshift distribution of galaxies. Thus, poorly localized GW events yield weaker constraints on cosmological parameters when compared to well-localized events.
@@
The latest LVK cosmological analyses [LIGOScientific:2021aug, LIGOScientific:2025jau] leverage GW data in combination with the set of galaxy catalogs called glade+ [Dalya:2018cnd, Dalya:2021ewn]. This set of catalogs is essentially empty in the K-band at $z\sim 0.3$. Since many GW signals occur at higher equivalent redshifts (equivalent assuming a reference cosmology to convert the GW luminosity distance to a redshift), the catalog incompleteness poses a bottleneck. To formulate a completeness correction (i.e. what fraction of galaxies is observed for a given redshift), a galaxy magnitude distribution is required, the parameters of which are fixed in all standard codes, although this distribution is not known perfectly. Additionally, the aforementioned codes [Mastrogiovanni:2023emh, Gray:2023wgj, Borghi:2023opd] implement the completion correction assuming a uniform-in-comoving-distance distribution for the GW host. While it is acknowledged that the missing galaxies are not uniformly scattered, limited progress has been made to complete galaxy catalogs disposing of this unphysical assumption. For methodological development for a full galaxy catalog-GW analysis see [2019arXiv191112337L], and refer to [Oguri:2016dgk, Bera:2020jhx, Mukherjee:2022afz, Ghosh:2023ksl, Begnoni:2024tvj, Afroz:2024joi, Ghosh:2024cwc, Ferri:2024amc, Pedrotti:2025tfg] for a vivid recent interest treating the cross-correlation of GWs with galaxies (both for current and future detector networks, forecasting $H_{0}$ measurements at percent-level precision). Finally the work of [Finke:2021aom] has introduced a phenomenological prescription to account for spatial correlation between galaxies, while a completion technique based on the variance of the galaxy number counts in three-dimensional volumes of similar completion has been put forward in [Dalang:2023ehp, Dalang:2024gfk].
@@
Our works solves the aforementioned problems. We address the incompleteness, by reconstructing the full galaxy field given a magnitude-limited catalog, under the hypothesis that the galaxy field respects a given two-point statistic (accounting for spatial correlation). We also marginalize over the galaxy magnitude distribution; it is jointly inferred along with the galaxy field. This leads to a more informative galaxy redshift-sky position prior for GW events that implies a more informative dark siren $H_{0}$ measurement, schematically illustrated in Sec. 1.2. While this work demonstrates the reconstruction on simulated data it prepares the method for its application to glade+.
@@ 1.2 Representative toy model
For simplicity, we linearize the relationship between redshift and luminosity distance, i.e. $d_{L}=c\,z/H_{0}$. All analyzed GW events are localized in the same sky pixel (i.e. element of the sky) that contains $\mathcal{O}(10^{3})$ galaxies (detected and non-detected). To illustrate the impact of the galaxy completeness we then vary the number of observed galaxies included in the catalog. We assume that the redshift uncertainty of each galaxy to follow a Gaussian distribution with standard deviation of $\sigma_{z}=0.01$. From the observed galaxy catalog one can build the line-of-sight redshift prior, the sum of the individual galaxy redshift probability density distributions. [^1] As pointed out in [Gair:2022zsa] this treatment is approximate, since one should consider a hierarchical model for the galaxy distribution. Currently however, this treatment is standard in all inference codes. Note that because the catalog is incomplete, the line-of-sight redshift prior has two contributions: the in-catalog part (informed by the observed galaxies) and out-of-catalog part (with the standard assumption that galaxies are distributed uniformly in comoving volume). We construct the in-catalog part as a sum of Gaussian distributions with their respective redshift error that modulate the volumetric $z^{2}$ prior. The out-of-catalog part is simply modeled as a $z^{2}$ distribution. If we define the sigmoid
@@
Figure 1: Toy model: schematic impact of the galaxy catalog completeness. Left panel: The redshift prior for one fiducial sky pixel, where different colors indicate different completeness. The catalogs transition to the homogeneous distribution at different redshifts: while the catalog of 10 galaxies (orange) reverts to the homogeneous distribution already at $z\sim 0.2$, the catalog of 800 galaxies (blue) only transitions at $z\sim 1$. This has important consequences on the $H_{0}$ posterior as the right panel illustrates: while the most incomplete catalog provides an almost uninformative $H_{0}$ posterior (orange), the more complete catalogs with 200 and 800 galaxies (green and blue), respectively, lead to a more informative $H_{0}$ measurement.
@@ 1.3 Novelty
Let us here clearly draw up the differences between our previous method of [Leyde:2024tov] and the present work. In our previous reconstruction we have estimated the galaxy number count per three-dimensional voxel. We had assumed a simplified setup, namely a cartesian box, a parametrized form for the galaxy magnitude distribution, and a local bias prescription linking dark matter (DM) density contrast and mean galaxy rate. For the DM model itself, we assumed a log-normal field with a two-point statistic that was estimated but of a parametrized form. Here, we go significantly further and
@@ 1.4 Paper structure
Figure 3: Schematic overview of the hierarchical inference problem. Starting from the true catalog, one identifies possible galaxies (step Object detection). We indicate the apparent magnitude (cf. Eq. (2.8)) of each galaxy with its plotted color (blue), as well as its size. Darker, larger circles represent brighter galaxy in apparent magnitude. In a second step (step Inference of galaxy properties) the parameters of each galaxy are inferred, independently for each object. Finally, the inference of the individual galaxy properties is improved by considering a joint prior on the overall distribution of galaxies (step Solve inverse problem). In this step, we start from the magnitude-limited catalog with the noisy galaxy properties and build possible configurations of the true catalog that are compatible with the observations. Let us stress here that uncertainties in the galaxy sky position are neglected throughout this analysis.
@@ 2 Forward-modeling of galaxy catalog data
$^{n_{{\rm c},\hat{I}}}$ Observed number count in bin $\hat{I}$ (measured quantities)
@@
Schematically, the likelihood can be written as $\mathbb{P}(\{\mathcal{D}\}|\Lambda)$, with $\{\mathcal{D}\}\coloneqq\{\mathcal{D}_{c}\}_{c\in C}$ the observed catalog where $c$ is a label running over all galaxies in the observed catalog $C$. The raw catalog data comes in the form of fluxes in different bands, or spectra from which the galaxy properties (redshift, composition, sky position, star formation rates, etc.) are deduced. We will refer to these individual properties as $\theta$. As such, $\theta$ is a $\mathcal{O}(3-20)$-dimensional vector. One possible way to characterize each galaxy is through posterior samples that approximate $\theta\sim\mathbb{P}(\theta|\mathcal{D})\,$, the individual posterior of the galaxy properties. The likelihood of observing the catalog data $\{\mathcal{D}\}$ given the hyperparameters $\Lambda$ can be computed through [^2] In this expression we omit terms relating to the selection effect and that depend on the overall rate. This equation is for illustrating purposes only: this approach would be computationally prohibitive and will not be used in the later analysis.
@@
marginalizing the joint prior of the true catalog, $\mathbb{P}(\{\theta_{c}\}_{c\in C}|\Lambda)$, over the individual galaxy properties. While future computational acceleration might allow the evaluation of the above integral, it is not the aim of this work to marginalize over the properties of $\mathcal{O}(10^{6-8})$ galaxies.
Therefore, we choose to represent the catalog data as the galaxies histogrammed over the sky, the redshift and the apparent magnitude. Since the redshift is uncertain, we bin each galaxy by its maximum-likelihood value in redshift. That is, our observable is a number count in each three-dimensional bin, which we denote as $^{n_{{\rm c},\hat{I}}}$, where $\hat{I}$ is the index in redshift, sky position and magnitude, and the hat indicating that the galaxies are binned via their noisy maximum-likelihood properties. In general, the maximum-likelihood values do not correspond to their true values, an effect accounted for in our modeling, as outlined below.
@@
$\mathbb{P}(\{\mathcal{D}\}|\Lambda)\approx\mathbb{P}(\{{n_{{\rm c},\hat{I}}}\}|\Lambda)\,.$ (2.2)
Two assumptions underlie the above approximation, namely (i) the probability of $\mathbb{P}(\mathcal{D}|\Lambda)$ is constant in each bin and, (ii) the observed data of each galaxy can be fully described by the bin in which it falls (here its maximum-likelihood value for its properties). While the first assumption can be satisfied by using a fine grid over the galaxy properties, the second assumption is stronger. If we assume the $\theta$ uncertainties to be known, this assumes the individual galaxy likelihood to be independent of other latent variables not included in the histogram, e.g. the time of observation, the type of galaxy, etc.
@@
One can consider the problem we attempt to solve as a hierarchical inference problem where the individually obtained galaxy properties (which were independently inferred from the other galaxy data) are improved through the formulation of a joint prior of galaxy properties of position and magnitude, accounting for the three-dimensional spatial correlation and magnitude distribution. This is illustrated by Fig. 3. Note that the posterior estimate of the joint galaxy properties generally differs from the estimate for the galaxy properties separately (which both in themselves will depend on the prior knowledge). Since the former estimate takes into account population-level knowledge of the galaxy catalog it tends to reduce the individual galaxy parameter uncertainties, an effect that is also known as shrinkage [2019arXiv191112337L].
@@ 2.2 Dark matter model
In order to reconstruct the full galaxy field, we choose to model the mean galaxy number in each three-dimensional voxel as a local function of the DM density contrast. [^3] As discussed later in Sec. 2.3, this is an approximation, as the real galaxy formation process is a complex function of the halo size, galaxy history, environment etc. To produce the latter (spatially correlated) field in spherical coordinates, we now describe the log-normal field [Coles:1991if] that models the DM density contrast. Since we produce DM realizations on a cartesian grid, but need to evaluate its density on a spherical grid, we detail our discretization scheme and how we interpolate from cartesian to spherical coordinates in App. B.
Following the methodology of [Leyde:2024tov] based on [Coles:1991if], we generate a log-normal field with a given power spectrum on a cartesian grid. The power spectrum is computed using a phenomenological model that we introduce below in Eq. (2.5), the parameters of which we fit jointly with the other remaining hyperparameters such as the DM density contrast. The log-normal field is generated by applying a transformation to a Gaussian random field, given by [Coles:1991if, Agrawal:2017khv]
@@
where we denote the Gaussian random field as $F_{\rm G}$, and $\sigma_{\text{G}}^{2}$ is the variance of the Gaussian random field (over many field realizations). The Gaussian field is first constructed by sampling from a multivariate normal distribution with a covariance matrix derived from the desired power spectrum, exploiting an analytical connection between power spectrum of the log-normal field and its underlying Gaussian random field [Coles:1991if]. The log-normal transformation of Eq. (2.4) ensures that the resulting field is bounded by 1-1 from below, as required for the DM density contrast. Within this formalism, we do not take into account the redshift dependence of the power spectrum. Upon the application to real data this will warrant additional attention. Upon varying the redshift over the range considered later (the redshift varying between $z=0.13$ and $z=0.91)$, the amplitude of power spectrum varies mostly by amplitude not shape, with the former changing by $\mathcal{O}(2-3)$ [Piras:2023aub].
@@
[…]
where $\vec{k}$ is the (three-dimensional) wave vector and $||.||$ is the Euclidean norm. We show two examples of the power spectrum and their corresponding log-normal fields in two dimensions in Fig. 4. Two reasons made us choose a phenomenological model for the DM power spectrum: (i) our previous work [Leyde:2024tov] found little impact of using a power spectrum that depends on cosmological parameters and (ii) it is faster to calculate.